Please answer this question for me.
A random sample of 100 students who took a statistics exam were asked their score on the exam. The mean score on the test was 50; the standard deviation was 10. The scores are normally distributed.
A) What percent of the students scored below 30?

B) What percent scored between 30 and 55?

Answer by stanbon(57342)

A random sample of 100 students who took a statistics exam were asked their score on the exam. The mean score on the test was 50; the standard deviation was 10. The scores are normally distributed.
A) What percent of the students scored below 30?
Find the z-score for 30:
z(30)=(30-50)/10 = -2
Find P(z<-2) on your z-chart
or use normalcdf(-10,-2) = 0.0228 on your
TI calculator.
----------------
B) What percent scored between 30 and 55?
Find the z=score for 55:
z(55)=(55-50)/10 = 1/2
Find P(-2 or use normalcdf(-2,0.5)= 0.6687
no your TI calculator.
============
Cheers,
Stan H.